Geraldine Consultants, Inc. is considering a project that has the following cash flows: Year Cash Flow 0 -$1,000 1 400 2 300 3 500 4 400 The company's WACC is 10%. What are the project's payback, internal rate of return, and net present value? Select one: a. Payback = 2.6, IRR = 21.22%, NPV = $300. b. Payback = 2.4, IRR = 21.22%, NPV = $260. c. Payback = 2.6, IRR = 24.12%, NPV = $300. d. Payback = 2.4, IRR = 10.00%, NPV = $600. e. Payback = 2.6, IRR = 21.22%, NPV = $260.
Answer #1
Year Cash flows Cumulative Cash flows
0 (1000) (1000)
1 400 (600)
2 300 (300)
3 500 200
4 400 600

Payback period=Last period with a negative cumulative cash flow+(Absolute value of cumulative cash flows at that period/Cash flow after that period).

=2+(300/500)

=2.6 years

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=400/1.0x+300/1.0x^2+500/1.0x^3+400/1.0x^4

Hence x=irr=21.22%(Approx).

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=400/1.1+300/1.1^2+500/1.1^3+400/1.1^4

=$1260(Approx)

NPV=Present value of inflows-Present value of outflows

=$1260-$1000

=$260(Approx)

Hence the correct option is:

e. Payback = 2.6, IRR = 21.22%, NPV = $260.